OpenStudy (anonymous):

(1+Xsquared+ysquared+xsquaredtimesysquared)dy=ysquareddx My mind is blank. I can't even figure out how to start this.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

So you have: \[\int\limits 1+x^2+y^2+xy^2 dy=\int\limits y^2 dx?\]. Or: \[(1+x^2+y^2+xy^2)dy=y^2 dx\].?

OpenStudy (anonymous):

Or some other form?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

xysquared should be xsquaredysquared

OpenStudy (anonymous):

Well 1+x^2+y^2+x^2y^2=(x^2+1)(y^2+1) So: (y^2+1)/y^2 dy=dx/(x^2+1) Integrating: \[\int\limits \frac{y^2+1}{y^2}=\int\limits \frac{dx}{x^2+1}\] \[\int\limits 1+y^{-2} dy=\int\limits \frac{dx}{x^2+1}\] \[y-y^{-1}=\arctan(x)\] Factoring out a y^(-1): \[\frac{y^2-1}{y}=\arctan(x)\].